OpenStudy (anonymous):

The relationship between the hours spent studying and the average grade on a science test has been proven to be linear. The average grade for a student who studied 1 hour was a 76% and for a student who studied 4 hours was a 94%. Write the linear equation representing this scenario and then find how many hours of studying would have resulted in a grade of 88%.

OpenStudy (anonymous):

o.o

OpenStudy (flying_dutchman):

1 hour = 76% | 4 hours = 94% There's a difference of 18% and 3 hours, so you can find the ratio easily. 1 = 76 2 = 82 3 = 88 4 = 94 Equation: Grade = 70 + 6h | h --> hours 88 = 70 + 6h h = (88 - 70) / 6 --> h=3

OpenStudy (anonymous):

Thank you for all that hard work. <3

OpenStudy (flying_dutchman):

you're welcome)